OpenStudy (smurfy14):

A bike accelerates uniformly from rest to a speed of 7.10 m/s over a distance of 25.4 m. Determine the acceleration of the bike.

OpenStudy (anonymous):

v^2=2as a= v^2/2a a= 7.1*7.1/25.4*2=0.998m/s^2

OpenStudy (anonymous):

-0.992 NEGATIVE

OpenStudy (smurfy14):

crap i meant 35.4 :/ but thanks anyways.